# 4M-21: An Any-to-Any Vision Model for Tens of Tasks and Modalities

### v1

@@ Abstract
Current multimodal and multitask foundation models like 4M [62] or UnifiedIO [59, 58] show promising results, but in practice their out-of-the-box abilities to accept diverse inputs and perform diverse tasks are limited by the (usually rather small) number of modalities and tasks they are trained on. In this paper, we expand upon the capabilities of them by training a single any-to-any model on tens of highly diverse modalities and by performing co-training on large-scale multimodal datasets and text corpora. This includes training on several semantic and geometric modalities, feature maps from recent state of the art models like DINOv2 and ImageBind, pseudo labels of specialist models like SAM and 4DHumans, and a range of new modalities that allow for novel ways to interact with the model and steer the generation, for example image metadata or color palettes. A crucial step in this process is performing discrete tokenization on various modalities, whether they are image-like, neural network feature maps, vectors, structured data like instance segmentation or human poses, or data that can be represented as text.
Through this, we expand on the out-of-the-box capabilities of multimodal models and specifically show the possibility of training one model to solve at least 3x more tasks/modalities than existing ones and doing so without a loss in performance. This enables more fine-grained and controllable multimodal generation capabilities and allows us to study the distillation of models trained on diverse data and objectives into one unified model. We successfully scale the training to a three billion parameter model using tens of modalities and different datasets. The resulting multimodal models and training code are open sourced at https://4m.epfl.ch.
Figure 1: We demonstrate the possibility of training a single model on tens of highly diverse modalities using a multimodal masking objective [62], without a loss in performance compared to existing specialized single/few task models trained to solve significantly fewer modalities. The modalities are mapped to discrete tokens using modality-specific tokenizers. The resulting model can generate any of the modalities from any subset of them.
11footnotetext: Equal contribution & corresponding authors. Randomized order.22footnotetext: Work partially done while at Apple or EPFL, respectively.
@@ 1 Introduction
Having a single neural network to handle a wide and varied range of tasks and modalities has been a longstanding goal. Such a model, especially when capable of any-to-any predictions, brings notable advantages, such as model size and test-time computational efficiency and enabling modality fusion.
However, multitask learning has commonly faced significant challenges. For example, the training often suffers from negative transfer and typically requires careful strategies for balancing losses or gradients [46, 93, 82, 94, 31]. Moreover, training a single network on tasks and modalities that vary greatly in terms of dimensionality, data type, and value ranges presents additional complexitiesModality vs task: “Modalities” usually denote the inputs to a model (e.g. sensory signals), and “tasks” usually denote the outputs (e.g. semantics). The adopted architecture in multimodal masked modeling enables a symmetric input-output structure, thus modalities and tasks are used interchangeably in this paper.. Recent notable efforts in the space of multimodal and multitask training, such as Pix2Seq [16, 17], OFA [88], Unified-IO [59, 58], or 4M [62], have made significant strides in unifying the representation space for conceptually different inputs and targets. A large part of their success can be attributed to transforming different modalities into a common representation, namely sequences of discrete tokens, and training relatively standard Transformer architectures on them. While these works show promising results, they are typically trained on a small set of modalities. This raises the question if increasing the set of tasks/modalities they can solve will lead to a degradation of performance.
@@
  • Modalities: Increase from 7 in the existing best any-to-any models to 21 diverse modalities, enabling new capabilities like cross-modal retrieval, controllable generation, and strong out-of-the-box performance. This is one of the first times in the vision community that a single model can solve tens of diverse tasks in an any-to-any manner (see fig. 2), without sacrificing performance and especially do so without any of the conventional multitask learning difficulties [73, 46, 93, 82, 94, 31].
@@
Figure 2: One-to-all generation. 4M-21 can generate all modalities from any given input modality using chained generation [62]. Notice the high consistency among the predictions of all modalities. Each row starts from a different modality coming from the same data sample. Highlighted in green are new input/output pairs that 4M [62] cannot predict nor accept as input. Note that, while this figure shows predictions from a single input, 4M-21 can generate any modality from any subset of all modalities.
@@ 3 Multimodal capabilities
Figure 4: Fine-grained & steerable multimodal generation. Top left: 4M-21 can generate variants of images that are grounded in any input modality, here human poses. Bottom left: This enables us to perform multimodal edits (e.g. editing the shape of a polygon or grounding generation with edges) and probe the learned representation. For example, by only changing the shape of the ellipse, 4M-21 renders the bowl from different angles. Top right: By pre-training on 21 types of modalities, including T5-XXL embeddings, and co-training on a large text corpus, we show improved text understanding capabilities (even when the input is captions instead of language model embeddings). Bottom right: Compared to generating images from captions only, metadata provides a more direct and steerable way of controlling the multimodal data generation process, enabling exciting further research into generative dataset design.
@@ 3.2 Multimodal retrieval
Table 1: Out-of-the-box (zero-shot) performance. We show the performance for a common subset of tasks: surface normals and depth estimation on DIODE [86], semantic and instance segmentation on COCO [55], k-NN retrieval on ImageNet-1K [75], and 3D human keypoint estimation on 3DPW [87]. We compare to a set of strong baselines and specialist models, including our pseudo labelers. Our model learned to solve all these tasks without loss of performance and is significantly better than the baselines and competitive with pseudo labelers, while being a single model for all tasks. ✗ denotes that a given model cannot solve the task out-of-the-box. * shows the tokenizer reconstruction quality and provides an estimate on the performance upper bound due to tokenization. See fig. 13 for qualitative comparisons. Best results are bolded, second best underlined.
@@ 3.3 Evaluating out-of-the-box capabilities
Our model consistently achieves strong out-of-the-box performance, and often matches or even outperforms the pseudo labelers and other specialist baselines, while being a single model for all tasks. Notice the large performance gap with other multitask models like Unified-IO [59] and Unified-IO-2 [58]. For kNN retrieval, 4M-21 XL performance approaches the tokenizer bound, i.e. the retrieval performance using the DINOv2 tokenizer reconstructions. While our smaller models lag behind 4M models, we observe that 4M-21 XL is able to match the performance of 4M-7 XL, but is capable of interfacing with three times the number of modalities. It is expected that additional model capacity is needed, as it is trained to solve many more tasks.
@@ 4 Transfer experiments
To study the scaling characteristics of pre-training any-to-any models on a much larger set of modalities, we train models across three different sizes: B, L, and XL. We then transfer their encoders to downstream tasks and evaluate on both unimodal (RGB) and multimodal (RGB + Depth) settings. The decoders are discarded for all transfer experiments, and we instead train task-specific heads. We perform self-comparisons in a similar manner to [62, 7], as well as comparing to a set of strong baselines.
@@
The results are shown in table 2. We make the following observations: 1) for the transfer tasks that are similar to the seven modalities of 4M, e.g. semantic segmentation or depth, 4M-21 does not lose performance due to being trained on many more modalities, 2) for novel transfer tasks like 3D object detection that are sufficiently different from 4M modalities, we observe an improved performance. Moreover, the performance improves with larger model sizes, showing promising scaling trends. These trends can be further seen in the multimodal transfer results, which will be explained next.
@@ 6 Limitations and Discussion
We demonstrate training an any-to-any model on tens of diverse modalities and tasks. This is achieved by mapping all modalities to discrete sets of tokens via modality-specific tokenizers and using a multimodal masked training objective [62]. We successfully scaled the training to three billion parameters and to 21 modalities and different datasets, without a degradation in performance compared to the existing more specialized single/few task models. This results in strong out-of-the-box capabilities as well new potential for multimodal interaction, generation, and retrieval, all by a single unified model. Below, we discuss limitations of our method and future work.
Transfer/emergent capabilities: One hope from multitask training is leading to a model that can solve novel tasks, often referred to as “transfer” or “emergent” capabilities. While a multitask model brings several key advantages even without transfer/emergence advantages (using a single model for broad out-of-the-box capabilities without sacrificing performance, modality fusion, etc.), and we showed success at them, we observe that the potential in transfer/emergence improvement remains largely untapped. In general, compared to LLMs, vision/multimodal models in the community have not shown exciting results in terms of transfer/emergence yet. We find this to be an important point for us and the community to address in the future, e.g., via designing multitask architectures that have emergence, in contrast to out-of-the-box capabilities, as their main objective.
Better tokenization: Like any token-based model, 4M-21 can directly benefit from progress on tokenizers, e.g. higher reconstruction fidelity.
Co-training on partially aligned datasets: We showed the possibility of training on partially aligned datasets, e.g. text data from C4 and other modalities from CC12M, yet further investigations and a larger mixture of datasets are expected to bring stronger capabilities, which we aim as future work.

### v2

@@ Abstract
Current multimodal and multitask foundation models, like 4M [62] or UnifiedIO [59, 58], show promising results. However, their out-of-the-box abilities to accept diverse inputs and perform diverse tasks are limited by the (usually small) number of modalities and tasks they are trained on. In this paper, we develop a single any-to-any model trained on tens of highly diverse modalities and by performing co-training on large-scale multimodal datasets and text corpora. This includes training on images and text along with several semantic and geometric modalities, feature maps from recent state of the art models like DINOv2 and ImageBind, pseudo labels of specialist models like SAM and 4DHumans, and a range of new modalities that allow for novel ways to interact with the model and steer the generation, for example, image metadata or color palettes. A crucial step in this process is performing discrete tokenization on various modalities, whether they are image-like, neural network feature maps, vectors, structured data like instance segmentation or human poses, or data that can be represented as text.
Through this, we show the possibility of training one model to solve at least 3x more tasks/modalities than existing models and doing so without a loss in performance. In addition, this enables more fine-grained and controllable multimodal generation capabilities and allows studying the distillation of models trained on diverse data and objectives into one unified model. We scale the training to a three billion parameter and different datasets. The multimodal models and training code are open sourced at https://4m.epfl.ch.
Figure 1: We demonstrate training a single model on tens of highly diverse modalities without a loss in performance compared to specialized single/few task models. The modalities are mapped to discrete tokens using modality-specific tokenizers. The model can generate any of the modalities from any subset of them.
11footnotetext: Equal contribution & corresponding authors. Randomized order.22footnotetext: Work partially done while at EPFL and Apple.
@@ 1 Introduction
Having a single neural network to handle a wide and varied range of tasks and modalities has been a longstanding goal. Such a model, especially when capable of any-to-any predictions, brings notable advantages, such as test-time computational efficiency, model size, and enabling modality fusion.
However, multitask learning has commonly faced significant challenges. For example, the training often suffers from negative transfer, leads to reduction in performance compared to single-task models, and typically requires careful strategies for balancing losses or gradients [46, 93, 82, 94, 31]. Moreover, training a single network on tasks and modalities that vary greatly in terms of dimensionality, data type, and value ranges presents additional complexitiesModality vs task: “Modalities” usually denote the inputs to a model (e.g. sensory signals), and “tasks” usually denote the outputs (e.g. semantics). The adopted architecture in multimodal masked modeling enables a symmetric input-output structure, thus modalities and tasks are used interchangeably in this paper.. Recent notable efforts in the space of multimodal and multitask training, such as Pix2Seq [16, 17], OFA [88], 4M [62], or Unified-IO [59, 58] have made significant strides in unifying the representation space for conceptually different inputs and targets. A large part of their success can be attributed to transforming different modalities into a common representation, namely sequences of discrete tokens, and training relatively standard Transformer architectures on them. While these works show promising results, they are typically trained on a small set of modalities. This raises the question if increasing the set of tasks/modalities the models can solve will lead to a degradation of performance.
@@
  • Modalities: Increase from 7 in the existing best any-to-any models [62] to 21 diverse modalities, enabling new capabilities like cross-modal retrieval, controllable generation, and strong out-of-the-box performance. This is one of the first times in the vision community that a single model can solve tens of diverse tasks in an any-to-any manner (see fig. 2), without sacrificing performance and especially do so without any of the conventional multitask learning difficulties [73, 46, 93, 82, 94, 31].
@@
Figure 2: One-to-all generation. 4M-21 can generate all modalities from any given input modality and can benefit from chained generation [62]. Notice the high consistency among the predictions of all modalities for one input. Each row starts from a different modality coming from the same scene. Highlighted in green are new input/output pairs that 4M [62] cannot predict nor accept as input. Note that, while this figure shows predictions from a single input, 4M-21 can generate any modality from any subset of all modalities.
@@ 3 Multimodal capabilities
Figure 4: Fine-grained & steerable multimodal generation. Top left: 4M-21 can generate variants of images that are grounded in any input modality, here human poses. Bottom left: This enables us to perform multimodal edits (e.g. editing the shape of a polygon or grounding generation with edges) and probe the learned representation. For example, by only changing the shape of the ellipse, 4M-21 renders the bowl from different angles. Top right: By pre-training on 21 types of modalities, including T5-XXL embeddings, and co-training with language modeling on a large text corpus, we show improved text understanding capabilities (even when the input is captions instead of language model embeddings). Bottom right: Compared to generating images from captions only, metadata provides a more direct and steerable way of controlling the multimodal data generation process, enabling exciting further research into generative dataset design.
@@ 3.2 Multimodal retrieval
Table 1: Out-of-the-box (zero-shot) performance. We show the performance for a common subset of tasks: surface normals and depth estimation on DIODE [86], semantic and instance segmentation on COCO [55], k-NN retrieval on ImageNet-1K [75], and 3D human keypoint estimation on 3DPW [87]. We compare to a set of strong baselines and specialist models, including our pseudo labelers. The model learned to solve all the tasks without a loss of performance, is significantly better than the baselines, and is competitive with pseudo labelers, while being a single model for all tasks. Compared to 4M-7, the 4M-21 model preserved its performance while solving 3x more tasks. ✗ denotes that a given model cannot solve the task out-of-the-box. * shows the tokenizer reconstruction quality and provides an estimate on the performance upper bound due to tokenization. See fig. 13 for qualitative comparisons. Best results are bolded, second best underlined.
@@ 3.3 Evaluating out-of-the-box capabilities
Our model consistently achieves strong out-of-the-box performance, and often matches or even outperforms the pseudo labelers and other specialist baselines, while being a single model for all tasks. Notice the large performance gap with other multitask models like Unified-IO [59] and Unified-IO-2 [58]. For kNN retrieval, 4M-21 XL performance approaches the tokenizer bound, i.e. the retrieval performance using the DINOv2 tokenizer reconstructions. While the smaller models lag behind 4M models, we observe that 4M-21 XL is able to match the performance of 4M-7 XL, while being trained to solve three times more tasks. The trend over the model size needing to be larger is expected as the number of tasks increase.
@@ 4 Transfer experiments
To study the scaling characteristics of pre-training any-to-any models on a larger set of modalities, we train models across three different sizes: B, L, and XL. We then transfer their encoders to downstream tasks and evaluate on both unimodal (RGB) and multimodal (RGB + Depth) settings. The decoders are discarded for all transfer experiments, and we instead train task-specific heads. We perform self-comparisons in a similar manner to [62, 7], as well as comparing to a set of strong baselines.
@@
The results are shown in table 2. We make the following observations: 1) for the transfer tasks that are similar to the seven modalities of 4M, e.g. semantic segmentation or depth, 4M-21 does not lose performance due to being trained on many more modalities, 2) for novel transfer tasks like 3D object detection that are sufficiently different from 4M modalities, we observe an improved performance. Moreover, the performance improves with larger model sizes, showing promising scaling trends. These trends can be further seen in the multimodal transfer results, which we explain next.
@@ 6 Limitations and Discussion
We developed an any-to-any model on tens of diverse modalities and tasks. This was achieved by mapping all modalities to discrete sets of tokens via modality-specific tokenizers and using a multimodal masked training objective [62]. We successfully scaled the training to 3 billion parameters and to 21 modalities and different datasets, without a degradation in performance compared to the existing expert single/few task models. This results in strong out-of-the-box capabilities as well new potential for multimodal interaction, generation, and retrieval, all by a single unified model. Below, we discuss limitations and future work.
Transfer/emergent capabilities: One hope from training a single network on several tasks is leading to a model that can solve novel tasks, often referred to as “transfer” or “emergent” capabilities. While, as we showed, a multitask model brings several key advantages even without transfer/emergence (e.g., efficiency, using a single model for broad out-of-the-box capabilities, modality fusion, etc.), we observe that the potential for transfer/emergence remains largely untapped. In general, compared to LLMs, vision/multimodal models have not shown exciting results in terms of transfer/emergence yet. We find this to be an important point to address in the future, e.g., via designing multitask architectures that have emergence, in contrast to out-of-the-box capabilities, as their main objective.
Better tokenization: Like any token-based model, ours can directly benefit from progress on tokenizers, e.g. higher reconstruction fidelity.
Co-training on partially aligned datasets: We showed the possibility of training on partially aligned datasets, e.g. text data from C4 and other modalities from CC12M, yet further investigations and a larger mixture of datasets are expected to bring stronger capabilities.